OpenStudy (anonymous):

estimate the difference first. then subtract. simplify, if necessary. the problem is 3 3/5 - 7/10 = thanks

OpenStudy (anonymous):

3 3/5 is mix number?

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

than 3 3/5 is 18/5

OpenStudy (anonymous):

\[\frac{18}{5}-\frac{7}{10}\]

OpenStudy (anonymous):

ok i have that so far.

OpenStudy (anonymous):

\[\frac{(18*2)-7}{10}\]

OpenStudy (anonymous):

\[\frac{36-7}{10}=\frac{29}{10}\]

OpenStudy (anonymous):

okay where did the 2 come from?

OpenStudy (anonymous):

denominator is 10 5*2=10

OpenStudy (anonymous):

never mind that the number for denominator right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay got you. i have it im writin it down.

OpenStudy (anonymous):

denominator is 10 5*2=10

OpenStudy (anonymous):

k got that far.

OpenStudy (anonymous):

solve the problem. 18 1/6- 7 1/2 could you take me through this

OpenStudy (anonymous):

\[18\frac{1}{6}-7\frac{1}{2}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

u know how to write fraction?

OpenStudy (anonymous):

no not on the computer

OpenStudy (anonymous):

but yes on paper

OpenStudy (anonymous):

no not here,i mean you know conver to fraction

OpenStudy (anonymous):

yes i do

OpenStudy (anonymous):

good

OpenStudy (anonymous):

\[\frac{109}{6}-\frac{15}{2}\]r u ok from here?

OpenStudy (anonymous):

where did you get the 15 over 2 from?

OpenStudy (anonymous):

7(1/2)=(7*2)+1=15 keep 2

OpenStudy (anonymous):

18(1/6)=(18*6)+1=109 keep 6

OpenStudy (anonymous):

it was not 2 its 7 1/12

OpenStudy (anonymous):

I ask u said yes

OpenStudy (anonymous):

on the bottom can u rewrite?

OpenStudy (anonymous):

im sorry i apologize.

OpenStudy (anonymous):

7(1/12) or 71/12

OpenStudy (anonymous):

18 1/6= 18(1/6)?

OpenStudy (anonymous):

yes thats right.

OpenStudy (anonymous):

7(1/12) ?

OpenStudy (anonymous):

okay would you mind helping me .

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I will rewrite made sure it correct equation

OpenStudy (anonymous):

solve for n 1. n+1 3/8=2 1/4 2. 6 2/3=n-8 1/6 n okay thats fine. i apologize that i don't know how to help you understand

OpenStudy (anonymous):

u need help the mix number?

OpenStudy (anonymous):

do the make it a fraction?

OpenStudy (anonymous):

\[18\frac{1}{6}-7\frac{1}{12}\]?

OpenStudy (anonymous):

here it the problem cause thats not it. solve for n 1. n+1 3/8=2 1/4 2. 6 2/3=n-8 1/6

OpenStudy (anonymous):

I still don't get

OpenStudy (anonymous):

n+1 time (3/8)=2 and 1/4?

OpenStudy (anonymous):

\[n+1\frac{3}{8}=2\frac{1}{4}\]?

OpenStudy (anonymous):

yes that is correct

OpenStudy (anonymous):

\[1\frac{3}{8}=(1*8)+3=\frac{11}{8}\]

OpenStudy (anonymous):

\[2\frac{1}{4}=(2*4)+1=\frac{9}{4}\]

OpenStudy (anonymous):

ok got that part so far

OpenStudy (anonymous):

\[n+\frac{11}{8}=\frac{9}{4}\]

OpenStudy (anonymous):

\[n=\frac{9}{4}-\frac{11}{8}\]

OpenStudy (anonymous):

denominator is 8, you need made 9/4 have the same 8 by time 2

OpenStudy (anonymous):

\[n=\frac{9*2}{4*2}-\frac{11}{8}=\frac{18}{8}-\frac{11}{8}\]

OpenStudy (anonymous):

k i got that so far

OpenStudy (anonymous):

\[n=\frac{7}{8}\]

OpenStudy (anonymous):

check : 7/8+11/8=18/8 = 9/4

OpenStudy (anonymous):

okay i have that. n when your done do you plug n chug

OpenStudy (anonymous):

and would u b able to help me with problem 2

OpenStudy (anonymous):

check:\[\frac{7}{8}+ \frac{11}{8}=\frac{18}{8}=\frac{9}{4}\]

OpenStudy (anonymous):

2. 6 2/3=n-8 1/6

OpenStudy (anonymous):

\[6\frac{2}{3}=n-8\frac{1}{6}\]?

OpenStudy (anonymous):

yes correct

OpenStudy (anonymous):

u know conver fraction right ?

OpenStudy (anonymous):

its been a while. no

OpenStudy (anonymous):

\[6\frac{2}{3}=(6*3)+2=20\]\[6\frac{2}{3}=\frac{20}{3}\]

OpenStudy (anonymous):

\[8\frac{1}{6}=(8*6)+1=49\rightarrow \frac{49}{6}\]

OpenStudy (anonymous):

\[n=\frac{20}{3}+\frac{49}{6}\] denominator is 6 keep 49/6 20/3 time 2 , then denominator is the same 6

OpenStudy (anonymous):

\[n=\frac{20*2}{3*2}+\frac{49}{6}\]

OpenStudy (anonymous):

\[n=\frac{40+49}{6}=\frac{89}{6}\]

OpenStudy (anonymous):

I did check it correct

OpenStudy (anonymous):

is the answer 14 n 5/6

OpenStudy (anonymous):

yes the same

OpenStudy (anonymous):

\[14\frac{5}{6}=(14*6)+5=89 \rightarrow = \frac{89}{6}\]is you onver to fraction is the same

OpenStudy (anonymous):

k yeah i understand

OpenStudy (anonymous):

good cluck your home work

OpenStudy (anonymous):

thanks i really appreciate it